OpenStudy (anonymous):

88 is 80% of what? please show work

OpenStudy (anonymous):

\[88\times \frac{100}{80}=110\]

OpenStudy (anonymous):

88/0.8 = 110; This is just the inverse of the above answer -- instead of multiplying by 100/80, simply divide by 80/100 (or 0.80)